# From Pixels to Insights: A Survey on Automatic Chart Understanding in the Era of Large Foundation Models

### v1

@@ Abstract
Data visualization in the form of charts plays a pivotal role in data analysis, offering critical insights and aiding in informed decision-making. Automatic chart understanding has witnessed significant advancements with the rise of large foundation models in recent years. Foundation models, such as generative pre-trained transformers (GPT), have revolutionized various natural language processing (NLP) tasks and are increasingly being applied to chart understanding tasks. This survey paper provides a comprehensive overview of the recent developments, challenges, and future directions in chart understanding within the context of these foundation models. The paper begins with a background section, defining chart understanding, outlining problem formulations, and discussing fundamental building blocks crucial for studying chart understanding tasks, including vision encoders, chart-to-table translation, OCR modules, text encoders, and text decoders. In the section on tasks and datasets, we explore various tasks within chart understanding, including chart question answering, chart captioning, chart-to-table conversion , chart fact-checking, and chart caption factual error correction. We discuss evaluation metrics and sources of both charts and textual inputs. Modeling strategies are then examined, encompassing both classification-based and generation-based approaches, along with tool augmentation techniques that enhance chart understanding performance. Furthermore, we discuss the state-of-the-art performance of each task and discuss how we can improve the performance. Challenges and future directions are addressed in a dedicated section, highlighting issues such as domain-specific charts, and evaluation criteria concerning faithfulness, coverage, relevancy, robustness, fairness, and data bias. Additionally, we delve into dissecting components of these multimodal foundation models for chart understanding tasks, including the necessity of tuning LM backbones, the efficacy of multi-stage training processes, and the potential sufficiency of synthetically generated data. Agent-oriented settings, where chart understanding systems interact with users or other systems, are also explored. Lastly, we discuss related tasks such as natural image understanding, table understanding, and document understanding, providing insights into the broader landscape of visual and textual data comprehension. This survey paper serves as a comprehensive resource for researchers and practitioners in the fields of natural language processing, computer vision, and data analysis, providing valuable insights and directions for future research in chart understanding leveraging large foundation models. The studies mentioned in this paper, along with emerging new research, will be continually updated at: https://github.com/khuangaf/Awesome-Chart-Understanding.
@@ 1 Introduction
The Significance of Chart Understanding in Information Communication In our contemporary world of multimedia information, where the volume and complexity of data continue to burgeon, the role of charts stands paramount in facilitating coherent and insightful communication of factual information, conveying insights, and making decisions. Across diverse fields spanning academia, scientific research, digital media, and business realms, charts serve as indispensable tools for translating raw data into comprehensible visual narratives. Their ability to encapsulate complex datasets in a concise and intuitive format empowers decision-makers to grasp key insights swiftly, aiding informed reasoning and strategic planning. Recognizing this pivotal role of charts in modern information dissemination, there has been a sustained interest within the computational community, as evidenced by a plethora of research in automatic chart understanding. In particular, works on chart question answering [1, 2, 3, 4, 5], chart captioning [6, 7, 8, 9], chart-to-table conversion [10, 11], chart fact-checking [12, 13], chart caption factual error correction [14] have laid foundational frameworks for exploring the intricacies of chart semantics in chart understanding technologies.
@@
Challenges and Opportunities in Chart Understanding Amidst the Era of Large Foundation Models  Traditional chart understanding work [15, 7] focuses on finetuning methods that generally experience limitations with respect to domain portability and reasoning robustness. Excitingly, the advent of large vision-language foundation models (e.g., GPT-4V [16], LLaVA [17]) has spawned a paradigm shift in automated reasoning capabilities, catalyzing unprecedented advancements across various multimedia cognitive tasks – including advancements in strong zero/few-shot reasoning capabilities through text-based prompting. But amidst this landscape of transformative innovation, the domain of chart understanding still remains entrenched in inherent complexity and formidable challenges. Charts present a unique set of obstacles owing to their multifaceted visual representations and nuanced semantics. From bar charts and line graphs to pie charts and scatter plots, each chart type employs a distinct visual syntax to convey data relationships, requiring sophisticated interpretative mechanisms beyond mere pixel-level pattern recognition. Charts serve as a conduit for unearthing profound insights such as emerging trends, outliers that challenge assumptions, and relationships among variables that may not be immediately apparent from raw data in tabular form alone. They enable comparative analyses across data points, offering a visual platform for juxtaposing varied entities or time periods succinctly. Furthermore, the intrinsic diversity of underlying datasets, ranging from simple numerical relations to intricate multidimensional entities, adds another layer of complexity to the chart understanding task. Despite these challenges, automated chart understanding sits at the intersection of opportunity and impact, offering a gateway to unlock actionable insights buried within the pixels of visual narratives. By harnessing the capabilities of large foundation models, chart understanding demonstrates improved potential in bridging the gap between raw visual data and meaningful insights, thereby enabling scalable technology for accessible applications and enhanced human cognition.
@@
Several studies have surveyed the landscape of chart understanding research, yet these surveys often manifest certain gaps in comprehensiveness or specificity. Some surveys do not cover modern datasets employed in chart understanding research, as well as the most contemporary modeling approaches, such as those involving pre-trained vision-language models and large foundation models [18, 19]. Conversely, other surveys concentrate predominantly on the visualization aspect (i.e. the transformation of data into charts), thereby sidelining the nuanced task of chart interpretation [20, 21, 22, 23]. This survey paper aims to bridge these gaps.
@@
We first define automatic chart understanding and the fundamental building blocks to the problem formulation in Section 2. We discuss the multifaceted nature of chart understanding, encompassing tasks from interpreting chart visuals to analyzing underlying data, as well as outline the structural modeling components integral to chart understanding, such as vision encoders, OCR modules, and text decoders, and their role in transforming raw chart images and text queries into meaningful insights. Then, in Section 3, we examine the datasets driving chart understanding research and the metrics for model evaluation. This section analyzes the sources, diversity, and limitations of these datasets, providing insights into the current chart understanding data landscape. It also reviews various evaluation metrics, underlining the necessity for robust and nuanced assessment methods. With insights from these characteristics, we further provide trending popular modeling strategies for automatic chart understanding. Section 4 delves into the diverse modeling strategies in chart understanding, including adaptations from natural image understanding, vision-language pre-training, and foundation models like large language models (LLMs) and large vision-language models (LVLMs). In particular, we emphasize the impact of choices in vision encoders and text decoders on model effectiveness and discuss tool augmentation’s role in chart understanding. We conclude this section by showcasing the state-of-the-art performance on different chart understanding tasks and how we can improve upon them.
@@
Finally, Section 5 addresses the challenges and future directions in chart understanding. We highlight the importance of domain-specific charts, the need for comprehensive evaluation metrics, and the potential for adversarial settings to enhance model robustness and versatility. We also discuss in Section 6 how chart understanding stands amidst the intersection of related works in natural image understanding [24], table understanding [25], and document understanding [26].This survey paper concludes by identifying key areas for future research, such as developing models for complex charts, refining evaluation metrics, and diversifying datasets. We not only offer an in-depth overview of the current state of chart understanding but also set the stage for future advancements in this exciting intersection of data visualization and machine learning.
@@ 2.2 Problem Formulation and Fundamental Building Blocks
TABLE I: Notations used in this paper.
Notation Description
$\mathbf{c}$
Input chart image
$\mathbf{H}_{c}$
Representation matrix for the input chart image
$\mathbf{q}$
Input textual query
$\mathbf{H}_{q}$
Representation matrix for the input textual query
${q}_{i}$
The $i$-th token of the input textual query.
$l^{q}$
Number of tokens in the input textual query
$\hat{\mathbf{s}}$
Scene text extracted from the input chart image
$\mathbf{y}$
Gold standard textual response
$\hat{\mathbf{y}}$
Predicted textual response
${\mathbf{T}}$
Gold standard underlying data table of the input chart image
${\mathbf{T}}_{i}$
The $i$-th token for the data table $\mathbf{T}$ after linearization
$\hat{\mathbf{T}}$
Data table extracted from the input chart image by a system
$\textsc{Encoder}_{\text{vis}}$
Vision encoder
$\textsc{Encoder}_{\text{txt}}$
Text encoder
$\textsc{Decoder}_{\text{txt}}$
Text decoder
$\textsc{Projector}_{\text{vt}}$
Vision-language decoder
$\mathcal{X}$
Input context set to the text decoder
@@ 3 Tasks and Datasets
Dataset Tasks Source Domain # Instances Real-world Data Real-world Charts FigureQA [2] FQA Synthetic General 2M DVQA [1] FQA Synthetic General 3M LEAF-QA [32] FQA World Development Indicators et al. General 2M LEAF-QA++ [33] FQA World Development Indicators et al. General 2M PlotQA [4] FQA/C2T World Bank Open Data et al. General 29M ChartQA [3] FQA/C2T Statista/Pew General 10K MapQA [34] FQA Kaiser Family Foundation General 796K PaperQA [35] FQA Scientific Papers Science 107 SciGraphQA [36] FQA Arxiv Science 296K MMC-Bench [37] FQA Statista et al. General 2K ChartBench [38] FQA Kaggle General 17K ArXivQA [39] FQA ArXiv Science 100K OpenCQA [5] OQA Pew General 8K FigCAP [40] CAP Synthetic General 101K Chart2text [41] CAP Statista General 8K SciCap [6] CAP ArXiv Science 476K LineCap [42] CAP ArXiv Science 4K ChaTa+ [43] CAP ArXiv/World Health Organization Science 2K Chart-to-Text [8] CAP/C2T Statista/Pew General 44K ChartSumm [44] CAP Statista/Knoema General 84K VisText [9] CAP/C2T Statista General 12K FigCaps-HF [45] CAP ArXiv Science 134K ArxivCap [39] CAP ArXiv Science 4M ChartFC [13] FC Wikipedia General 16K ChartCheck [12] FC Wikimedia General 10K Chocolate [14] FEC/FID Statista/Pew General 1K ✓ / ✗
@@ 4.2.3 Large Vision-language Models
Model # Instances # Tasks Real-world Data Real-world Charts Synthetic Data Augmentation Pre-training Datasets ChartT5 [59] 495K 2 ✓ / ✗ ✓ / ✗ MatCha [37] 46M 5 ✓ / ✗ ✓ / ✗ UniChart [37] 7M 4 ✓ / ✗ ✓ / ✗ Instruction-tuning Datasets MMCA [37] 2M 9 ✓ / ✗ ChartLlama [27] 160K 7 ChartAssistant [28] 39M 5 ✓ / ✗ ChartInstruct [66] 191K 9
@@ 5.2 Evaluation
Faithfulness. The faithfulness of a model’s output refers to the extent to which the information presented is accurate and consistent with the data depicted in the chart. Unfaithfulness or factual inconsistency can lead to outputs that misinform users, potentially causing misinterpretations or incorrect decision-making based on inaccurate chart understanding. This is of particular significance in domains where accurate data representation is critical, such as education and news reporting. Although the ChartVE metric [14] offers a robust assessment of the factual consistency between chart data and textual outputs, it is unable to perform granular evaluations that locate errors within text spans. Future studies can emphasize the development of finer-grained evaluation techniques alongside the classification of error types, such as the value errors and trend errors delineated by previous work [14], or the confidence boundary of model output hallucinations [102]. Such approaches would provide a clearer picture of where models are failing in terms of faithfulness.
@@ 5.5 Multilingual Chart Understanding
6 Related Tasks
6.1 Natural Image Understanding
6.2 Table Understanding
Table understanding involves interpreting information presented in tabular data. This task aims to recognize the table’s layout and structure, identify the relationships between its rows and columns, and comprehend the semantic context embedded within the table’s content. In contrast to charts, which employ graphical elements to visually represent data, tables arrange this data in a meticulously organized format. They utilize rows and columns to systematically display information, offering a direct view of the data’s structure and interrelationships.
6.3 Document Understanding

### v2

@@ Abstract
Data visualization in the form of charts plays a pivotal role in data analysis, offering critical insights and aiding in informed decision-making. Automatic chart understanding has witnessed significant advancements with the rise of large foundation models in recent years. Foundation models, such as large language models, have revolutionized various natural language processing tasks and are increasingly being applied to chart understanding tasks. This survey paper provides a comprehensive overview of the recent developments, challenges, and future directions in chart understanding within the context of these foundation models. We review fundamental building blocks crucial for studying chart understanding tasks. Additionally, we explore various tasks and their evaluation metrics and sources of both charts and textual inputs. Various modeling strategies are then examined, encompassing both classification-based and generation-based approaches, along with tool augmentation techniques that enhance chart understanding performance. Furthermore, we discuss the state-of-the-art performance of each task and discuss how we can improve the performance. Challenges and future directions are addressed, highlighting the importance of several topics, such as domain-specific charts, lack of efforts in developing evaluation metrics, and agent-oriented settings. This survey paper serves as a comprehensive resource for researchers and practitioners in the fields of natural language processing, computer vision, and data analysis, providing valuable insights and directions for future research in chart understanding leveraging large foundation models. The studies mentioned in this paper, along with emerging new research, will be continually updated at: https://github.com/khuangaf/Awesome-Chart-Understanding.
@@ 1 Introduction
The Significance of Chart Understanding in Information Communication In our contemporary world of multimedia information, where the volume and complexity of data continue to burgeon, the role of charts stands paramount in facilitating effective communication of factual information, conveying insights, and informing decision making. Across diverse fields spanning academia, scientific research, digital media, and business realms, charts serve as indispensable tools for translating raw data into comprehensible visual narratives. Their ability to encapsulate complex datasets in a concise and intuitive format empowers decision-makers to grasp key insights swiftly, aiding informed reasoning and strategic planning. Recognizing this pivotal role of charts in modern information dissemination, there has been a sustained interest within the computational community, as evidenced by a plethora of research in automatic chart understanding. In particular, works on chart question answering [1, 2, 3, 4, 5], chart captioning [6, 7, 8, 9], chart-to-table conversion [10, 11], chart fact-checking [12, 13], chart caption factual error correction [14] have laid foundational frameworks for exploring the intricacies of chart semantics in chart understanding technologies.
@@
Challenges and Opportunities in Chart Understanding Amidst the Era of Large Foundation Models  Traditional chart understanding work [15, 7] focuses on finetuning methods that generally experience limitations with respect to domain portability and reasoning robustness. Excitingly, the advent of large vision-language foundation models (e.g., GPT-4V [16], LLaVA [17]) has spawned a paradigm shift in automated reasoning capabilities, catalyzing unprecedented advances across various multimedia cognitive tasks – including progress in strong zero/few-shot reasoning capabilities through text-based prompting. Despite the impressive improvements, the domain of chart understanding still presents major challenges. Charts present a unique set of obstacles owing to their rich visual representations and nuanced semantics, delivered through complex and creative organization of information manifested in text, numerical data, plots, and graphics objects. From bar charts and line graphs to pie charts and scatter plots, each chart type employs a distinct visual syntax to convey data relationships, requiring sophisticated interpretative mechanisms beyond mere pixel-level pattern recognition. Charts serve as a conduit for unearthing profound insights such as emerging trends, outliers that challenge assumptions, and relationships among variables that may not be immediately apparent from raw data in tabular form alone. They enable comparative analyses across data points, offering a visual platform for juxtaposing varied entities or time periods succinctly. Furthermore, the intrinsic diversity of underlying datasets, ranging from simple numerical relations to intricate multidimensional entities, adds another layer of complexity to the chart understanding task. Despite these challenges, automated chart understanding offers a gateway to unlock actionable insights buried within the pixels of visual narratives. By harnessing the capabilities of large foundation models, chart understanding demonstrates improved potential in bridging the gap between raw visual data and meaningful insights, thereby enabling high-level information extraction, logical reasoning, and decision-making.
@@
Several studies have surveyed the landscape of chart understanding research, yet these surveys often show limitations in covering comprehensive scopes or reviewing information of sufficient specificity. Some surveys do not cover modern datasets employed in chart understanding research, as well as the most contemporary modeling approaches, such as those involving pre-trained vision-language models and large foundation models [18, 19]. Conversely, other surveys concentrate predominantly on the visualization aspect (i.e. the transformation of data into charts), thereby missing the nuanced task of chart interpretation [20, 21, 22, 23]. This survey paper aims to bridge these gaps.
@@
We first define automatic chart understanding and the fundamental building blocks to the problem formulation in Section 2. We discuss the multifaceted nature of chart understanding, encompassing tasks from interpreting chart visuals to analyzing underlying data. We also illustrate how chart understanding stands amidst the intersection of related works in natural image understanding [24], table understanding [25], and document understanding [26], as well as outling the structural modeling components integral to chart understanding, such as vision encoders, OCR modules, and text decoders, and their role in transforming raw chart images and text queries into meaningful insights. Then, in Section 3, we examine the datasets driving chart understanding research and the metrics for model evaluation. This section analyzes the sources, diversity, and limitations of these datasets, providing insights into the current chart understanding data landscape. It also reviews various evaluation metrics, underlining the necessity for robust and nuanced assessment methods. With insights from these characteristics, we further provide trending popular modeling strategies for automatic chart understanding. Section 4 delves into the diverse modeling strategies in chart understanding, including adaptations from natural image understanding, vision-language pre-training, and foundation models like large language models (LLMs) and large vision-language models (LVLMs). In particular, we emphasize the impact of choices in vision encoders and text decoders on model effectiveness and discuss tool augmentation’s role in chart understanding. We conclude this section by showcasing the state-of-the-art performance on different chart understanding tasks and how we can improve upon them.
@@
Finally, Section 5 addresses the challenges and future directions in chart understanding. We highlight the importance of domain-specific charts, the need for comprehensive evaluation metrics, and the potential for adversarial settings to enhance model robustness and versatility. This survey paper concludes by identifying key areas for future research, such as developing models for complex charts, refining evaluation metrics, and diversifying datasets. We not only offer an in-depth overview of the current state of chart understanding but also set the stage for future advancements in this exciting intersection of data visualization and machine learning.
@@ 2.1 What is Automatic Chart Understanding?
2.2 Related Tasks
2.2.1 Natural Image Understanding
2.2.2 Table Understanding
Table understanding involves interpreting information presented in tabular data [31, 32, 33, 34]. This task aims to recognize the table’s layout and structure, identify the relationships between its rows and columns, and comprehend the semantic context embedded within the table’s content. In contrast to charts, which employ graphical elements to visually represent data, tables arrange this data in a meticulously organized format. They utilize rows and columns to systematically display information, offering a direct view of the data’s structure and interrelationships.
2.2.3 Document Understanding
@@ 2.3 Problem Formulation and Fundamental Building Blocks
TABLE I: Notations used in this paper.
Notation Description
$\mathbf{c}$ Input chart image
$\mathbf{H}_{c}$ Representation matrix for the input chart image
$\mathbf{q}$ Input textual query
$\mathbf{H}_{q}$ Representation matrix for the input textual query
${q}_{i}$ The $i$-th token of the input textual query.
$l^{q}$ Number of tokens in the input textual query
$\hat{\mathbf{s}}$ Scene text extracted from the input chart image
$\mathbf{y}$ Gold standard textual response
$\hat{\mathbf{y}}$ Predicted textual response
${\mathbf{T}}$ Gold standard underlying data table of the input chart image
${\mathbf{T}}_{i}$ The $i$-th token for the data table $\mathbf{T}$ after linearization
$\hat{\mathbf{T}}$ Data table extracted from the input chart image by a system
$\textsc{Encoder}_{\text{vis}}$ Vision encoder
$\textsc{Encoder}_{\text{txt}}$ Text encoder
$\textsc{Decoder}_{\text{txt}}$ Text decoder
$\textsc{Projector}_{\text{vt}}$ Vision-language decoder
$\mathcal{X}$ Input context set to the text decoder
@@ 3 Tasks and Datasets
Dataset Tasks Source Domain # Instances Real-world Data Real-world Charts FigureQA [2] FQA Synthetic General 2M DVQA [1] FQA Synthetic General 3M LEAF-QA [43] FQA World Development Indicators et al. General 2M LEAF-QA++ [44] FQA World Development Indicators et al. General 2M PlotQA [4] FQA/C2T World Bank Open Data et al. General 29M ChartQA [3] FQA/C2T Statista/Pew General 10K MapQA [45] FQA Kaiser Family Foundation General 796K PaperQA [46] FQA Scientific Papers Science 107 SciGraphQA [47] FQA Arxiv Science 296K MMC-Bench [48] FQA Statista et al. General 2K ChartBench [49] FQA Kaggle General 17K ArXivQA [50] FQA ArXiv Science 100K OpenCQA [5] OQA Pew General 8K FigCAP [51] CAP Synthetic General 101K Chart2text [52] CAP Statista General 8K SciCap [6] CAP ArXiv Science 476K LineCap [53] CAP ArXiv Science 4K ChaTa+ [54] CAP ArXiv/World Health Organization Science 2K Chart-to-Text [8] CAP/C2T Statista/Pew General 44K ChartSumm [55] CAP Statista/Knoema General 84K VisText [9] CAP/C2T Statista General 12K FigCaps-HF [56] CAP ArXiv Science 134K ArxivCap [50] CAP ArXiv Science 4M ChartFC [13] FC Wikipedia General 16K ChartCheck [12] FC Wikimedia General 10K Chocolate [14] FEC/FID Statista/Pew General 1K  /
@@ 4.2.3 Large Vision-language Models
Model # Instances # Tasks Real-world Data Real-world Charts Synthetic Data Augmentation Pre-training Datasets ChartT5 [70] 495K 2  /  / MatCha [48] 46M 5  /  / UniChart [48] 7M 4  /  / Instruction-tuning Datasets MMCA [48] 2M 9  / ChartLlama [38] 160K 7 ChartAssistant [39] 39M 5  / ChartInstruct [77] 191K 9
@@ 5.2 Evaluation
Faithfulness. The faithfulness of a model’s output refers to the extent to which the information presented is accurate and consistent with the data depicted in the chart. Unfaithfulness or factual inconsistency can lead to outputs that misinform users [113, 114, 115], potentially causing misinterpretations or incorrect decision-making based on inaccurate chart understanding. This is of particular significance in domains where accurate data representation is critical, such as education and news reporting. Although the ChartVE metric [14] offers a robust assessment of the factual consistency between chart data and textual outputs, it is unable to perform granular evaluations that locate errors within text spans. Future studies can emphasize the development of finer-grained evaluation techniques alongside the classification of error types, such as the value errors and trend errors delineated by previous work [14], or the confidence boundary of model output hallucinations [116]. Such approaches would provide a clearer picture of where models are failing in terms of faithfulness.